OpenStudy (anonymous):

choose the ordered pair (x,y) that is a solution of the equation. -12x+4y=-4 answers; (2,-12) (-1,-9) (-2,-12) (-3,-10)

OpenStudy (anonymous):

@texaschic101 (;

OpenStudy (texaschic101):

-12x + 4y = -4 (2,-12) -12(2) + 4(-12) = -4 -24 - 48 = -4 not this one (-1,-9) -12(-1) + 4(-9) = -4 12 - 36 = -4 not this one (-2,-12) -12(-2) + 4(-12) = -4 24 - 48 = -4 not this one (-3,-10) -12(-3) + 4(-10) = -4 36 - 40 = -4 -4 = -4 (correct) It would have to be the last one...lol